# VLN-MME : Diagnosing MLLMs as Language-guided Visual Navigation agents

### v1

@@ 4.1 Settings
GPT-5 9.72 6.41 36.50 54.00 28.19 41.02 18.09 33.03 9.91 8.26 25.50 42.50 20.73 35.06 11.04 29.07 10.02 4.62 45.50 69.50 23.65 31.09 14.05 25.08
Gemini-2.5 Pro 9.57 6.28 32.00 49.50 24.31 41.28 18.16 33.21 9.77 8.14 24.00 41.50 16.88 35.27 11.13 29.25 9.87 4.54 41.00 64.00 19.94 31.24 14.17 25.23
InternVL3-2B 2.37 8.55 4.00 6.00 3.26 32.83 2.97 33.54 2.45 9.58 3.67 5.00 3.31 27.97 2.81 31.14 9.72 5.85 25.00 44.00 4.50 15.52 3.04 17.39
InternVL3-8B 5.85 7.80 16.50 28.50 10.89 36.05 11.48 35.18 6.49 8.80 12.67 19.33 6.64 27.19 6.91 28.24 6.23 5.50 30.00 40.50 10.36 26.92 8.12 25.74
LLaVA-OV-7B 8.35 8.50 10.00 20.00 5.50 26.62 5.50 23.33 8.20 9.47 11.00 19.00 6.00 25.84 6.00 25.16 9.46 6.11 15.50 33.50 2.02 15.90 2.53 15.57
Qwen2.5-VL-7B 10.41 7.12 26.50 41.00 10.12 27.97 12.88 25.38 9.60 8.67 15.67 23.67 6.00 23.29 7.62 20.63 10.15 4.90 33.50 46.00 7.41 17.18 6.33 17.34
@@
GPT-5 9.82 6.59 34.50 52.00 26.13 42.07 16.04 34.09 10.03 8.39 24.50 40.50 18.67 36.05 9.03 30.02 10.11 4.76 42.50 66.50 21.79 33.08 12.07 27.03
Gemini-2.5 Pro 9.66 6.46 30.00 47.50 22.28 42.26 16.11 34.24 9.88 8.27 22.00 39.50 15.41 36.28 9.16 30.17 9.96 4.68 38.00 61.00 17.86 33.27 12.12 27.16
InternVL3-2B 7.52 8.82 9.00 20.00 4.88 23.45 4.98 24.08 7.35 9.95 4.67 12.67 1.61 19.60 2.02 21.94 7.26 6.15 18.00 35.50 4.74 22.73 3.77 22.14
InternVL3-8B 8.54 8.42 18.00 34.50 10.84 28.16 10.84 29.42 9.64 9.76 13.00 24.00 6.06 20.36 6.75 22.16 8.95 5.38 33.50 53.50 8.01 16.44 5.85 19.04
LLaVA-OV-7B 8.35 8.50 13.00 23.00 5.81 25.55 6.04 25.47 8.14 9.56 9.00 17.67 4.53 21.74 4.86 21.98 8.53 5.74 24.50 37.00 5.45 21.01 4.35 20.49
Qwen2.5-VL-7B 6.29 9.07 14.00 20.50 7.12 25.82 7.03 25.41 7.09 8.94 10.67 18.67 6.68 25.66 7.27 25.81 7.45 5.90 25.50 36.50 6.78 22.67 5.23 21.56
@@
We evaluate two proprietary and four open-source MLLMs in a zero-shot setting: GPT5, Gemini2.5 Pro, Qwen2.5-VL-7B (bai2025qwen2), InternVL3-2/8B (zhu2025internvl3), LLaVA-One-Vision-7B (li2024llava). These models are integrated into eight distinct agent configurations, categorized into two primary classes: agents using text summarization and text map as memory. Each class includes four variants: a baseline, one with CoT prompting, one with reflection-based reasoning, and one featuring both CoT and reflection. All open-source MLLMs are served using the vLLM backend (kwon2023efficient) to ensure efficient inference and memory management. We assess their performance on all the tasks in our benchmark, additionally, we compare these zero-shot agents against previously finetuned VLN specialist and finetuned MLLM agents on the R2R and REVERIE tasks, evaluating performance across both the full dataset from prior evaluation methods and our benchmark. All experiments are conducted on a single NVIDIA A100 GPU with 40GB VRAM.
@@ 4.2 Performance Analysis
Counterintuitively, the integration of advanced prompting strategies like CoT or reflection does not consistently yield performance improvements and often proves detrimental. As shown in Table 2, applying CoT to Qwen2.5-VL-7B in the Fine-Grained task drops the SR from 27.5% to 21%. Similarly, the distinction between Text Summarization and Text Map memory architectures is less critical than expected. Performance variances are generally minimal, although specific smaller models, such as InternVL3-2B, derive a tangible benefit from structured map memory in Coarse-grained tasks, improving from 7.33% SR to 12% SR. We argue that these counterintuitive limitations stem from the models’ deficient embodied context awareness. Our investigation reveals two critical, interrelated issues. First, MLLMs exhibit a strong bias toward “local reasoning,” relying heavily on immediate visual inputs and current instructions while neglecting the rich historical context. Second, due to this lack of historical grounding, the agents struggle to anticipate the downstream consequences of their actions, failing to adapt or recover from errors within the long-term sequential flow of the task. The comprehensive error analysis in Section 4.3 provides further evidence supporting this conclusion.
@@ 4.3 Discussion
As discussed in section 4.2, we reveal some counterfactual behavior when MLLMs performing embodied navigation. We further conduct an error analysis to understand their error pattern and find that they are hindered by fundamental limitations across several cognitive dimensions (see more details in Appendix E. Interestingly, we find that the high navigation failure rate is overwhelmingly dominated by looping behaviors, shown in Figure 4. It is not a superficial issue but symptomatic of deeper challenges in instruction fidelity, spatial reasoning, historical context utilization, and the grounding of multi-modal perception into action. We discuss these three interconnected aspects below.
@@ 4.4 Diagnostic Case Study
+ Oracle Assistant (Qwen3VL) 52.00 68.00 34.00
@@
1-shot Failure Example 12.00 24.00 9.00
2-shot Failure Examples 16.00 28.00 11.00
3-shot Failure Examples 16.00 32.00 14.00
@@ 5 Conclusion
In this work, we investigate MLLMs as zero-shot agents using Vision-and-Language Navigation as a challenging downstream task to evaluate core agentic capabilities. We introduce VLN-MME, a unified, modular, and simulator-free evaluation framework designed to probe MLLMs’ spatial reasoning, understanding of action consequences, and long-horizon sequential decision making, rather than to establish a new VLN benchmark. Our results show that current MLLMs exhibit severe limitations in translating perception into action, leading to poor zero-shot performance. Through fine-grained, step-level error analysis, VLN-MME reveals systematic failure patterns and provides concrete guidance for chain-of-thought–oriented data generation and post-training of embodied vision-language models.
@@ E.1 Trajectory Visualization and Analysis Tool
This section details the custom tool developed for the qualitative analysis of agent trajectories. The tool, named the VLN Result Visualizer, developing using Gradio, provides an interactive interface for a step-by-step inspection of any navigation episode, which is crucial for understanding the nuances of agent behavior beyond aggregate metrics.

### v2

@@ 4.1 Settings
GPT-5 9.72 6.41 41.02 18.09 33.03 54.00 36.50 28.19 9.91 8.26 35.06 11.04 29.07 42.50 25.50 20.73 10.02 4.62 31.09 14.05 25.08 69.50 45.50 23.65
Gemini-2.5 Pro 9.57 6.28 41.28 18.16 33.21 49.50 32.00 24.31 9.77 8.14 35.27 11.13 29.25 41.50 24.00 16.88 9.87 4.54 31.24 14.17 25.23 64.00 41.00 19.94
InternVL3-2B 2.37 8.55 32.83 2.97 33.54 6.00 4.00 3.26 2.45 9.58 27.97 2.81 31.14 5.00 3.67 3.31 9.72 5.85 15.52 3.04 17.39 44.00 25.00 4.50
InternVL3-8B 5.85 7.80 36.05 11.48 35.18 28.50 16.50 10.89 6.49 8.80 27.19 6.91 28.24 19.33 12.67 6.64 6.23 5.50 26.92 8.12 25.74 40.50 30.00 10.36
LLaVA-OV-7B 8.35 8.50 26.62 5.50 23.33 20.00 10.00 5.50 8.20 9.47 25.84 6.00 25.16 19.00 11.00 6.00 9.46 6.11 15.90 2.53 15.57 33.50 15.50 2.02
Qwen2.5-VL-7B 10.41 7.12 27.97 12.88 25.38 41.00 26.50 10.12 9.60 8.67 23.29 7.62 20.63 23.67 15.67 6.00 10.15 4.90 17.18 6.33 17.34 46.00 33.50 7.41
@@
GPT-5 9.82 6.59 42.07 16.04 34.09 52.00 34.50 26.13 10.03 8.39 36.05 9.03 30.02 40.50 24.50 18.67 10.11 4.76 33.08 12.07 27.03 66.50 42.50 21.79
Gemini-2.5 Pro 9.66 6.46 42.26 16.11 34.24 47.50 30.00 22.28 9.88 8.27 36.28 9.16 30.17 39.50 22.00 15.41 9.96 4.68 33.27 12.12 27.16 61.00 38.00 17.86
InternVL3-2B 7.52 8.82 23.45 4.98 24.08 20.00 9.00 4.88 7.35 9.95 19.60 2.02 21.94 12.67 4.67 1.61 7.26 6.15 22.73 3.77 22.14 35.50 18.00 4.74
InternVL3-8B 8.54 8.42 28.16 10.84 29.42 34.50 18.00 10.84 9.64 9.76 20.36 6.75 22.16 24.00 13.00 6.06 8.95 5.38 16.44 5.85 19.04 53.50 33.50 8.01
LLaVA-OV-7B 8.35 8.50 25.55 6.04 25.47 23.00 13.00 5.81 8.14 9.56 21.74 4.86 21.98 17.67 9.00 4.53 8.53 5.74 21.01 4.35 20.49 37.00 24.50 5.45
Qwen2.5-VL-7B 6.29 9.07 25.82 7.03 25.41 20.50 14.00 7.12 7.09 8.94 25.66 7.27 25.81 18.67 10.67 6.68 7.45 5.90 22.67 5.23 21.56 36.50 25.50 6.78
@@
We evaluate two proprietary and four open-source MLLMs in a zero-shot setting: GPT5, Gemini2.5 Pro, Qwen2.5-VL-7B (bai2025qwen2), InternVL3-2/8B (zhu2025internvl3), LLaVA-One-Vision-7B (li2024llava). These models are integrated into eight distinct agent configurations, categorized into two primary classes: agents using text summarization and text map as memory. Each class includes four variants: a baseline, one with CoT prompting, one with reflection-based reasoning, and one featuring both CoT and reflection. All open-source MLLMs are served using the vLLM backend (kwon2023efficient) to ensure efficient inference and memory management. We assess their performance on all the tasks in our benchmark. Additionally, we compare these zero-shot agents against previously finetuned VLN specialist and finetuned MLLM agents on the R2R and REVERIE tasks, evaluating performance across both the full dataset from prior evaluation methods and our benchmark. All experiments are conducted on a single NVIDIA A100 GPU with 40GB VRAM.
@@ 4.2 Performance Analysis
Counterintuitively, the integration of advanced prompting strategies like CoT or reflection does not consistently yield performance improvements and often proves detrimental. As shown in Table 2, applying CoT to Qwen2.5-VL-7B in the Fine-Grained task drops the SR from 27.5% to 21%. Similarly, the distinction between Text Summarization and Text Map memory architectures is less critical than expected. Performance variances are generally minimal, although specific smaller models, such as InternVL3-2B, derive a tangible benefit from structured map memory in Coarse-grained tasks, improving from 7.33% SR to 12% SR. We argue that these counterintuitive limitations stem from the models’ deficient embodied context awareness. Our investigation reveals two critical, interrelated issues. First, MLLMs exhibit a strong bias toward “local reasoning”, relying heavily on immediate visual inputs and current instructions while neglecting the rich historical context. Second, due to this lack of historical grounding, the agents struggle to anticipate the downstream consequences of their actions, failing to adapt or recover from errors within the long-term sequential flow of the task. The comprehensive error analysis in Section 4.3 provides further evidence supporting this conclusion.
@@ 4.3 Discussion
As discussed in section 4.2, we reveal some counterfactual behavior when MLLMs perform embodied navigation. We further conduct an error analysis to understand their error pattern and find that they are hindered by fundamental limitations across several cognitive dimensions (see more details in Appendix E). Interestingly, we find that the high navigation failure rate is overwhelmingly dominated by looping behaviors, shown in Figure 4. It is not a superficial issue but symptomatic of deeper challenges in instruction fidelity, spatial reasoning, historical context utilization, and the grounding of multi-modal perception into action. We discuss these three interconnected aspects below.
@@ 4.4 Diagnostic Case Study
+ Oracle Assistant (Qwen3VL) 52.00 68.00 41.28
@@
1-shot Failure Example 12.00 24.00 9.42
2-shot Failure Examples 16.00 28.00 11.25
3-shot Failure Examples 16.00 32.00 14.29
@@ 5 Conclusion
In this work, we introduce VLN-MME, a modular, simulator-free framework for diagnosing MLLMs as zero-shot embodied agents. Our evaluation reveals a critical "perception-action gap": models possess robust visual grounding yet fail at sequential decision-making, with Chain-of-Thought strategies often proving detrimental due to limited context awareness. Crucially, our diagnostic study on hard negatives confirms that these failures stem primarily from deficient strategic planning rather than perceptual limitations, as evidenced by substantial performance recovery under oracle guidance. Ultimately, VLN-MME highlights that future MLLM development must prioritize bridging the gap between static visual understanding and dynamic, history-aware strategic reasoning.
Limitation
We present and discuss the limitations of our work to outline the scope of our analysis. First, due to computational and temporal constraints, our evaluation covers a representative selection of MLLMs and agent designs rather than an exhaustive spectrum. However, we mitigate this by releasing a highly modular interface that enables the community to seamlessly integrate and benchmark emerging models. Second, while our framework supports broader capabilities such as dialog-based navigation and multi-language instructions, we restricted our current analysis to standard instruction following to establish a rigorous baseline. We designed the dataset construction pipeline to be extensible, allowing future researchers to easily incorporate these diverse tasks. Finally, this work is primarily diagnostic: we identify critical deficiencies in MLLM 3D spatial reasoning and strategic planning but do not propose specific algorithmic remedies. We posit these findings as a strategic guide for future work in embodied MLLM post-training and agent design.
@@ E.1 Trajectory Visualization and Analysis Tool
This section details the custom tool developed for the qualitative analysis of agent trajectories. The tool, named the VLN Result Visualizer, developed using Gradio, provides an interactive interface for a step-by-step inspection of any navigation episode, which is crucial for understanding the nuances of agent behavior beyond aggregate metrics.